OpenStudy (katielong):

answer in standard form: sqrt(1.4x10^8)

OpenStudy (katielong):

\[\sqrt{1.4\times10^8}\]

OpenStudy (anonymous):

Is that an 8 or a 3?

OpenStudy (katielong):

8 sorry

OpenStudy (anonymous):

\[\sqrt{1.4 \times 100000000}\]

OpenStudy (anonymous):

\[\sqrt{140000000}\]

OpenStudy (anonymous):

\[2000 \sqrt{35}\]

OpenStudy (anonymous):

would be the final answer

OpenStudy (katielong):

it should be in standard form though

OpenStudy (anonymous):

Hmm. Then It would have to be 11832.16

OpenStudy (katielong):

yeah, i tried thhat and it's wrong :L

OpenStudy (anonymous):

Did you round?

OpenStudy (anonymous):

The ENTIRE decimal answer would be 11832.1595662.

OpenStudy (katielong):

oh hand on its 1.2x10^4

OpenStudy (anonymous):

Yes, could be.